# kNN-CLIP: Retrieval Enables Training-Free Segmentation on Continually Expanding Large Vocabularies

### v1

@@ 3.2 Inference Using the Continually Expanding Embedding Databases
The embedding database established serves as the retrieval source during inference to enhance the capabilities for various downstream visual models, as shown in Figure 3. Given a test image along with a set of query mask proposals $M_{q}\in\mathcal{R}^{k\times h\times w}$ and corresponding feature embeddings acquired through the same method outlined in Section 3.1, we conduct a k-Nearest Neighbor (kNN) search for each query $e_{q}$ within the embedding database $D_{vec}$ as follows:
$\text{kNN}_{k}(e_{q})=\left\{(c_{j},s_{j}(\cdot,\cdot))\,|\,(e_{j},c_{j})\in D_{vec}\right\}_{j=1}^{k},s_{j}(e_{q},e_{j})=\frac{e_{q}\cdot e_{j}}{\|e_{q}\|\|e_{j}\|}$ (2)
@@
where we accumulate the class-wise similarity over the class vector and normalize it into a probability distribution through a softmax operation. We consider the higher similarity between instance embeddings with a more frequent retrieval reflects higher confidence in the kNN-based prediction. We then leverage the prediction logits to augment the raw prediction logits $P_{base}\in\mathcal{R}^{C}$ from a pre-trained visual model through a class-wise logit modification as follows:
@@ 4.1 Implementation Details
Database Construction. Our database is created by extracting features from each dataset’s training sets using DINOv2 (Oquab et al., 2023) with a ViT-Giant architecture, featuring 4 register tokens. We select "keys" as our feature representation. We resize the images to $518\times 518$ and acquire a image feature of dimensions $1536\times 37\times 37$, as the patch size of chosen ViT is $14$. We then conduct mask average pooling to shrink the feature dimension to $1536$. Then, we store the shrunken feature with dimension of $1536$ and its corresponding label $c$ into the database using FAISS. Note that we do not store past images alongside.
k-Nearest Neighbors Search. We employ FAISS for feature retrieval using cosine similarity metric. Our method uses a brute force approach for identifying the closest embeddings, but we also explore approximate search methods like Hierarchical Navigable Small Worlds (HNSW) to improve efficiency at inference time (Prabhu et al., 2023a). We shall compare HNSW approximation with exact-search and detail the tradeoffs in our ablations. The number of nearest neighbors retrieved is 16 across dense visual tasks, with further details in the ablation section.
Augmenting FC-CLIP. We integrate our methodology into the current state-of-the-art open-vocabulary segmentation method FC-CLIP, which generates query masks and assigns class probabilities for each mask, following Mask2Former (Cheng et al., 2022). We use a linear combination of the original predicted class probabilities, with the retrieved prediction class probabilities from the kNN-CLIP method. We subsequently follow FC-CLIP to fuse the in-vocabulary classification scores and the retrieval-augmented out-of-vocabulary classification scores via geometric ensemble. This allows us to enhance the out-of-distribution classification without losing pretraining knowledge.
@@ 4.2 Catastrophic Forgetting Restricts Open-Vocabulary Capabilities of Models
Setup. We adopt a class-incremental continual learning setting adopted from CoMFormer(Cermelli et al., 2023). We benchmark this on semantic segmentation benchmark of COCO Panoptic as base dataset (having 330K images and 80 classes) and ADE20K (Zhou et al., 2019) as continual dataset (having 27K images and 150 classes). We first train the models like FC-CLIP on COCO Panoptic and then incrementally update this base model on ADE20K, with 5, 10 and 30 classes per timestep and a computational budget of one epoch (Prabhu et al., 2023b; a). We follow Mask2Former (Cheng et al., 2022) to use their loss function on an FC-CLIP (Yu et al., 2023) model, and evaluate mIoU across all classes in COCO Panoptic and ADE20K respectively after continual training.
Results. As shown in Table 1, when using naïve continual learning baseline, regardless of the number of incremental class, we observe consistent performance degradation of around 6 mIoU and 24 mIoU on the continual learning dataset ADE20K and the base dataset COCO-Panoptic respectively. The degradation on ADE20K can be attributable to catastrophic forgetting when learning across timesteps. Moreover, this causes concept forgetting, i.e., degradation of the open-vocabulary performance by a large margin, as measured by performance on COCO-Panoptic. This indicates that there is an pressing need for techniques to allow segmentation models to continually expand their vocabulary across novel concepts without losing their open-vocabulary segmentation capabilities. We show that our proposed training-free method kNN-CLIP allows effective alleviation of concept forgetting leading to significant performance improvement of +7.2 mIoU in ADE20K and +1.8 mIoU in COCO Panoptic.
@@ 4.3 Comparison with Continual Segmentation Approaches
Setup. We compare our method against the popular supervised method, Mask2Former (Cheng et al., 2022) and current state-of-the-art continual learning method, CoMFormer (Cermelli et al., 2023). We use the same experimental protocol as previously described for our method. These supervised models were initially trained on the first 100 classes and subsequently updated incrementally with the next 50 classes. We then evaluated the performance of these models using the mean Intersection over Union (mIoU) metric across all 150 classes from the ADE20K dataset.
Results. As shown in Table 2, Mask2Former (Cheng et al., 2022) and CoMFormer (Cermelli et al., 2023), exhibit large performance degradation when we reduce the number of incremental classes within 50 classes. In contrast, kNN-CLIP demonstrates a substantial advantage in preserving previous knowledge and mitigating knowledge loss, and it is inherently unaffected by the number of incremental classes. This resilience is attributed to our dynamic and continuously expanding embedding database, which significantly enhances our method’s effectiveness in continual learning scenarios.
@@ 4.4 Retrieval Enhances Panoptic Segmentation
Setup. Our exploration extends to panoptic segmentation, further validating the discernability on instance-level open-vocabulary recognition. We evaluate our method on ADE20K and COCO Panoptic datasets. To assess the performance comprehensively, we employ three critical metrics: Panoptic Quality (PQ), Average Precision (AP), and mean Intersection over Union (mIoU).
@@
Improving Base Dataset Performance. We highlight that we observed improvements of +0.4 in PQ, +0.2 in AP, and +1.8 in mIoU for COCO Panoptic dataset itself. This highlights that using retrieval even from the baseline model’s training dataset, in this case, COCO Panoptic, significantly enhances segmentation accuracy. Our method complements advances in open-vocabulary panoptic segmentation tasks, with these results demonstrating the consistent improvement achieved.
Table 4: Open-vocabulary semantic segmentation performance. Retrieval augmentation demonstrates state-of-the-art performances on open-vocabulary semantic segmentation.
@@
LSeg+ (Li et al., 2022a) COCO Stuff 3.8 7.8 18.0 46.5 -
@@ 4.5 Retrieval Enhances Semantic Segmentation
Setup. Our study extends studying the impact of training-free continual vocabulary expansion of kNN-CLIP to semantic segmentation, testing it’s efficacy across dense prediction tasks. We analyze the performance of kNN-CLIP over five diverse datasets: ADE20K and A847 (Zhou et al., 2019), containing 27K images, Pascal Context(PC)-59/459 (Mottaghi et al., 2014), and Pascal VOC-21 (Everingham et al., 2010) containing 10K images. These datasets include pixel-level annotation on various granularity covering wide-range of semantic concepts from 150, 847, 21, 59, and 459 classes, respectively. We use the same FC-CLIP backbone as in Section 4.1. For all benchmarks, we leverage mIoU metric to evaluate segmentation performance.
@@
Table 6: Confidence Threshold and Confidence Weighting Ablations. Comparisons of different choices of hyper-parameters on A-847.
@@
Table 7: Number of Retrieved Nearest Neighbors Ablations. Comparisons of different choices of number of retrieved nearest neighbors on A-847 with $T=0.7$ and $\lambda=1.2$.
@@ 4.6 Comparisons with Retrieval-based Approaches
Comparison. We also compare our results with Hummingbird (Balazevic et al., 2024), which proposes a novel pretraining framework that is suitable for fast downstream adaptions. We compare our semantic segmentation results with Hummingbird’s best performing model on Pascal VOC and ADE20K. The results of both methods are derived by performing nearest neighbors retrieval from the same evaluated downstream datasets without any further training.
Results. As shown in Table 5, our method outperforms the best Hummingbird model with an increase of +8.0 and +5.5 mIoU in PASCAL VOC and ADE20K respectively. The result further showcases the strong performance and fast adaptability of our straightforward method over other retrieval-based approaches.
@@ 4.7 Ablations
Effect of Confidence Threshold and Confidence Weightings. We examine the impact of varying hyper-parameter settings on performance, focusing on two key parameters: the confidence threshold and the retrieved confidence weightings. The selection of the confidence threshold is primarily influenced by the accuracy of the kNN retrieval module across different datasets. Our experiments are carried out on the A-847 dataset. As shown in Table 6, the optimal performance, measured in terms of the highest mIoU, is achieved when the confidence threshold, denoted as $T$ in Section 3.2, is set to $0.7$, and the confidence weighting, represented by $\lambda$ in the same section, is set at $1.2$.
Effect of Numbers of Neighbors Retrieved. In dense visual tasks, such as semantic segmentation, our approach is to enhance the pseudo classification logits for each query mask by incorporating information from the 16 closest neighbors. Additionally, we conduct experiments to examine the impact of varying the number of nearest neighbors retrieved. The results are shown in Table 7. These experiments are carried out using the previously determined optimal hyper-parameters, where $T$ is set to $0.7$ and $\lambda$ is set to $1.2$.
Approximate Nearest Neighbors Search. To optimize our inference time, we explore approximate nearest neighbors search algorithms to reduce runtime. We apply Hierarchical Navigable Small World to our feature retrieval module. As expected, we observe a trade-off between inference time and model performance. As in Table 8, it can be seen that adopting HNSW leads to a faster runtime but makes compromises with mIoU. With medium size of support set, such as ADE20K including approximately 220k features, brute force search is preferred; if the size of support set is large, as for COCO, which has 860k features, approximate nearest neighbors search offers better balance between inference time and accuracy.
Inference Time. We assess the inference time based on our implementation of FC-CLIP, taking into account only the time required to infer an image. The retrieval database is preconstructed, ensuring that our evaluation focuses solely on the inference cost. By integrating an additional feature extractor, DINOv2, into the existing architecture, we observe a compromise in inference speed as shown in Table 8. Furthermore, the kNN search module contributes to the slower inference time due to our utilization of brute force search. We observe that inference speed can be enhanced by implementing approximate neighbor search techniques as alternatives to brute force search, but approximate neighbor search leads to a slightly worse performance.
@@ 5 Conclusion
In our study, we examine the constraints related to the vocabulary used in earlier open-vocabulary methods that rely on CLIP, and we also highlight how performance diminishes when Vision-Language Models (VLMs) are fine-tuned using downstream annotations. Instead of adopting continual learning approaches, we introduce a novel method called kNN-CLIP, which does not require training and utilizes nearest neighbors retrieval to predict dense visual tasks. To accommodate ever-growing vocabularies, we employ a dynamic retrieval database, circumventing the need for retraining. Our approach demonstrates strong performance across several datasets and showcases significant improvements over other methods.
Acknowledgements
This work was supported by the UKRI grant: Turing AI Fellowship EP/W002981/1 and EPSRC/MURI grant: EP/N019474/1. We thank the Royal Academy of Engineering and FiveAI for their support. AP is funded by Meta AI Grant No. DFR05540.

### v2

@@ 3.2 Inference Using the Continually Expanding Embedding Databases
The embedding database established serves as the retrieval source during inference to enhance the capabilities for various downstream visual models, as shown in Fig 3. Given a test image along with a set of query mask proposals $M_{q}\in\mathcal{R}^{k\times h\times w}$ and corresponding feature embeddings acquired using DINOv2 (Oquab et al., 2023) through the same method outlined in Sec 3.1, we conduct a k-Nearest neighbour (kNN) search for each query $e_{q}$ within the embedding database $D_{vec}$ as follows:
$\text{kNN}_{k}(e_{q})=\left\{(c_{j},s_{j}(\cdot,\cdot))\,|\,(e_{j},c_{j})\in D_{vec},s_{j}\in\mathbb{s}_{sorted}[:k]\right\}_{j=1}^{k},s_{j}(e_{q},e_{j})=\frac{e_{q}\cdot e_{j}}{\|e_{q}\|\|e_{j}\|}$ (2)
@@
where we accumulate the class-wise similarity over the class vector and normalize it into a probability distribution through a softmax operation. We consider the higher similarity between instance embeddings with a more frequent retrieval reflects higher confidence in the kNN-based prediction. We then leverage the prediction logits to augment the raw prediction logits $P_{base}\in\mathcal{R}^{C}$ from CLIP through a class-wise logit modification as follows:
@@
Figure 4: Augmenting FC-CLIP. We integrate the retrieval augmentation module to the state-of-the-art segmentation model, FC-CLIP. FC-CLIP includes an in-vocabulary branch and an out-of-vocabulary branch. We don’t shown the original out-of-vocabulary branch here for simplicity. We use kNN-CLIP to augment the out-of-vocabulary branch using DINOv2 features and retrieved information.
3.3 Augmenting FC-CLIP
We integrate our methodology into the current state-of-the-art open-vocabulary segmentation model, FC-CLIP (Yu et al., 2023), which contains three main components. The model includes a mask generator, an in-vocabulary classifier, and an out-of-vocabulary classifier. All these components are built on top of a shared CLIP backbone. The pixel decoder and mask decoder generate class-agnostic masks, following Mask2Former (Cheng et al., 2022). Then, by mask-pooling the pixel features from the pixel decoder using the class-agnostic masks proposed, the in-vocabulary classifier gives the in-domain class embeddings; by mask-pooling the CLIP features, the out-vocabulary classifier yields its class embeddings. The results from both classifiers are then fused to give the final prediction.
We use our retrieval augmentation module to enhance the performance of the out-of-vocabulary classifier of FC-CLIP, which is the main reason for the open-vocabulary capability of FC-CLIP. The original out-of-vocabulary classifier leverages the cross-modal CLIP model for open-vocabulary classification. As in Fig 4, to enable retrieval augmentation, we additionally pass the input image into a frozen pretrained DINOv2. We then conduct mask-pooling on the DINO features and obtain our query embeddings as in Sec 3.1. We then compute the retrieval augmented class probabilities as described in Sec 3.2. We fuse the retrieval augmented class probabilities with the original out-of-vocabulary class probabilities derived from CLIP linearly as in Eq. 4. Then, the retrieval augmented results of the out-of-vocabulary classifier are combined with the results of the in-vocabulary classifier to give a final class prediction, following FC-CLIP. We then derive the segmentation masks from mask predictions and class predictions following Mask2Former and FC-CLIP.
@@ 4.1 Implementation Details
Database Construction. Our database is created by extracting features from each dataset’s training sets using DINOv2 (Oquab et al., 2023) with a ViT-Giant architecture, featuring 4 register tokens (Darcet et al., 2023). We select "keys" as our feature representation. We resize the images to $518\times 518$ and acquire an image feature of dimensions $1536\times 37\times 37$, as the patch size of chosen ViT is $14$. We then conduct mask average pooling to shrink the feature dimension to $1536$. Then, we store the shrunken feature with a dimension of $1536$ and its corresponding label $c$ into the database using FAISS. On average, a stored feature occupies 6kB of space. Note that we do not store past images alongside.
k-Nearest neighbours Search. We employ FAISS for feature retrieval using the cosine similarity metric. Our method uses a brute force approach for identifying the closest embeddings, but we also explore approximate search methods like Hierarchical Navigable Small Worlds (HNSW) to improve efficiency at inference time (Prabhu et al., 2023a). We shall compare HNSW approximation with exact-search and detail the tradeoffs in our ablations. The number of nearest neighbours retrieved is 16 across dense visual tasks, with further details in the ablation section.
@@ 4.2 Catastrophic Forgetting Restricts Open-Vocabulary Capabilities of Models
Setup. We adopt a class-incremental continual learning setting adopted from CoMFormer(Cermelli et al., 2023). We benchmark this on the semantic segmentation benchmark of COCO Panoptic as a base dataset (having 330K images and 80 classes) and ADE20K (Zhou et al., 2019) as a continual dataset (having 27K images and 150 classes). We start from the model weights of FC-CLIP, which was trained on COCO Panoptic, and then incrementally update the base model on ADE20K, with 5, 10 and, 30 classes per timestep and a computational budget of one epoch (Prabhu et al., 2023b; a). We follow Mask2Former (Cheng et al., 2022) to use their loss function on an FC-CLIP (Yu et al., 2023) model and evaluate mIoU across all classes in COCO Panoptic and ADE20K respectively after continual training.
Results. As shown in Table 1, when using naïve continual learning baseline, regardless of the number of incremental classes, we observe consistent performance degradation of around 6 mIoU and 24 mIoU on the continual learning dataset ADE20K and the base dataset COCO-Panoptic respectively. The degradation on ADE20K can be attributable to catastrophic forgetting when learning across timesteps. Moreover, this causes concept forgetting, i.e., degradation of the open-vocabulary performance by a large margin, as measured by performance on COCO-Panoptic. This indicates that there is a pressing need for techniques to allow segmentation models to continually expand their vocabulary across novel concepts without losing their open-vocabulary segmentation capabilities. We show that our proposed training-free method kNN-CLIP allows effective alleviation of concept forgetting leading to significant performance improvement of +7.2 mIoU in ADE20K and +1.8 mIoU in COCO Panoptic.
@@ 4.3 Comparison with Continual Segmentation Approaches
Setup. We compare our method under the setting of continual learning against the popular supervised method, Mask2Former (Cheng et al., 2022) and the current state-of-the-art continual learning method, CoMFormer (Cermelli et al., 2023). We use the same experimental protocol as previously described for our method. These supervised models were initially trained on the first 100 classes and subsequently updated incrementally with the next 50 classes. We then evaluated the performance of these models using the mean Intersection over Union (mIoU) metric across all 150 classes from the ADE20K dataset.
Results. As shown in Table 2, Mask2Former (Cheng et al., 2022) and CoMFormer (Cermelli et al., 2023) exhibit large performance degradation when we reduce the number of incremental classes within 50 classes. In contrast, kNN-CLIP demonstrates a substantial advantage in preserving previous knowledge and mitigating knowledge loss, and it is inherently unaffected by the number of incremental classes. This resilience is attributed to our dynamic and continuously expanding embedding database, which significantly enhances our method’s effectiveness in continual learning scenarios.
@@ 4.4 Retrieval Enhances Panoptic Segmentation
Setup. Our exploration extends to panoptic segmentation, further validating the discernability of instance-level open-vocabulary recognition. We evaluate our method on ADE20K and COCO Panoptic datasets. To assess the performance comprehensively, we employ three critical metrics: Panoptic Quality (PQ), Average Precision (AP), and mean Intersection over Union (mIoU).
@@
Improving Base Dataset Performance. We highlight that we observed improvements of +0.4 in PQ, +0.2 in AP, and +1.8 in mIoU for COCO Panoptic dataset itself. This highlights that using retrieval even from the baseline model’s training dataset, in this case, COCO Panoptic, significantly enhances segmentation accuracy. Our method complements advances in open-vocabulary panoptic segmentation tasks, with these results demonstrating the consistent improvements achieved.
Table 4: Open-vocabulary semantic segmentation performance. Retrieval augmentation demonstrates state-of-the-art performances on open-vocabulary semantic segmentation. [^3] Our method aims to continually enhance the performance of open-vocabulary segmentation models. For evaluation, we use the training set of each dataset, which is unseen by other methods, to construct our embedding database. Our approach significantly improves the performance of zero-shot baselines in a continual learning setting.
@@
LSeg+ (Li et al., 2022) COCO Stuff 3.8 7.8 18.0 46.5 -
@@ 4.5 Retrieval Enhances Semantic Segmentation
Setup. Our study extends studying the impact of training-free continual vocabulary expansion of kNN-CLIP to semantic segmentation, testing its efficacy across dense prediction tasks. We analyze the performance of kNN-CLIP over five diverse datasets: ADE20K and A847 (Zhou et al., 2019), containing 27K images, Pascal Context(PC)-59/459 (Mottaghi et al., 2014), and Pascal VOC-21 (Everingham et al., 2010) containing 10K images. These datasets include pixel-level annotation on various granularity covering a wide-range of semantic concepts from 150, 847, 21, 59, and 459 classes, respectively. We use the same FC-CLIP backbone as in Sec 4.1. For all benchmarks, we leverage the mIoU metric to evaluate segmentation performance.
@@
Table 6: (Left) Confidence Threshold and Confidence Weighting Ablations. Comparisons of different choices of hyper-parameters on A-847. (Right) Number of Retrieved Nearest Neighbors Ablations. Comparisons of different choices of number of retrieved nearest neighbors on A-847 with $T=0.7$ and $\lambda=1.2$.
@@ 4.6 Comparisons with Retrieval-based Approaches
Comparison. We also compare our results with Hummingbird (Balazevic et al., 2023), which proposes a novel pretraining framework that is suitable for fast downstream adaptions. Hummingbird employs nearest neighbour retrieval for predicting the classes of patches in target images, while we use kNN to augment the prediction of each proposed class agnostic mask. We compare our semantic segmentation results with Hummingbird’s best-performing model on Pascal VOC and ADE20K. The results of both methods are derived by performing nearest neighbour retrieval from the same evaluated downstream datasets without any further training.
Results. As shown in Table 5, our method outperforms the best Hummingbird model with an increase of +8.0 and +5.5 mIoU in PASCAL VOC and ADE20K respectively. The result demonstrates that predictions based on class agnostic masks are more robust, and the information in image patches might not be sufficient for nearest neighbour retrieval. The result further showcases the strong performance and fast adaptability of our straightforward method over other retrieval-based approaches.
@@ 4.7 Ablations
Effect of Confidence Threshold and Confidence Weightings. We examine the impact of varying hyper-parameter settings on performance, focusing on two key parameters: the confidence threshold and the retrieved confidence weightings. The selection of the confidence threshold is primarily influenced by the accuracy of the kNN retrieval module across different datasets. Our experiments are carried out on the A-847 dataset. As shown in Table 6 (Left), the optimal performance, measured in terms of the highest mIoU, is achieved when the confidence threshold, denoted as $T$ in Sec 3.2, is set to $0.7$, and the confidence weighting, represented by $\lambda$ in the same section, is set at $1.2$.
Effect of Numbers of neighbours Retrieved. In dense visual tasks, such as semantic segmentation, our approach is to enhance the pseudo-logits for classification of each query mask by incorporating information from the 16 closest neighbours. Additionally, we conduct experiments to examine the impact of varying the number of nearest neighbours retrieved. The results are shown in Table 6 (Right). These experiments are carried out using the previously determined optimal hyper-parameters, where $T$ is set to $0.7$ and $\lambda$ is set to $1.2$.
Approximate Nearest Neighbour Search. To optimize our inference time, we explore approximate nearest neighbour search algorithms to reduce runtime. We apply Hierarchical Navigable Small World to our feature retrieval module. As expected, we observe a trade-off between inference time and model performance. As in Table 7, it can be seen that adopting HNSW leads to a faster runtime but makes compromises with mIoU. With the medium size of the support set, such as ADE20K including approximately 220k features, brute force search is preferred; if the size of the support set is large, as for COCO, which has 860k features, approximate nearest neighbour search offers a better balance between inference time and accuracy.
Inference Time. We assess the inference time based on our implementation of FC-CLIP, taking into account only the time required to infer an image. The retrieval database is preconstructed, ensuring that our evaluation focuses solely on the inference cost. By integrating an additional feature extractor, DINOv2, into the existing architecture, we observe a compromise in inference speed as shown in Table 7. Furthermore, the kNN search module contributes to the slower inference time due to our utilization of brute force search. We observe that inference speed can be enhanced by implementing approximate neighbour search techniques as alternatives to brute force search, but approximate neighbour search leads to a slightly worse performance.
@@ 5 Conclusion
In our study, we examine the constraints related to the vocabulary used in earlier open-vocabulary methods that rely on CLIP, and we also highlight how performance diminishes when Vision-Language Models (VLMs) are fine-tuned using downstream annotations. Instead of adopting continual learning approaches, we introduce a novel method called kNN-CLIP, which does not require training and utilizes nearest neighbour retrieval to predict dense visual tasks. To accommodate ever-growing vocabularies, we employ a dynamic retrieval database, circumventing the need for retraining. Our approach demonstrates strong performance across several datasets and showcases significant improvements over other methods.
Broader Impact. We aim to enable dense prediction models across a large and expanding vocabulary of classes. However, significant advancements are required before deployment in critical applications like self-driving cars or medical imaging, pending thorough testing and validation. Additionally, to prevent misuse, we explicitly prohibit our approach’s use in military or surveillance contexts.